this is my problem
P = A((1+r)-n
P= 20,000(1+ 0.06)-20
P = 22,000(1.06)-20 this is my question do i divide or multiply here
to get my answer

Answer by lwsshak3(11628)

this is my problem
P = A((1+r)-n
P= 20,000(1+ 0.06)-20
P = 22,000(1.06)-20
***
The formula is not written correctly for compound interest.
It should be written as follows:
A=P(1+r)^n, P=initial investment, r=interest rate per period, n=number of periods, A=amount after n periods.